3) Given: 3Cu(s) + 8HNO3(aq) à 3Cu(NO3)2(aq) + 2NO(g) + 4H2O(aq) Step 1 is done!
How many grams of HNO3 are needed to dissolve 11.45g of Cu?
Step 2: Mass to moles
Step 3: Mole to Mole (Ratio)
Step 4: Mole to Mass
1
Expert's answer
2020-12-01T14:22:23-0500

3Cu(s) + 8HNO3(aq) = 3Cu(NO3)2(aq) + 2NO(g) + 4H2O(aq)

n=m/M

M (Сu) = 63.5 g/mol

M (HNO3) = 63.01 g/mol

n (Cu)=11.45/63.5 = 0.18 mol

n (HNO3) = 8/3 x n (Cu) = 8/3 x 0.18 = 0.48 mol

m (HNO3) = n x M = 0.48 x 63.01 = 30.3 g